# Divided We Fall: Defending Against Adversarial Attacks via Soft-Gated Fractional Mixture-of-Experts with Randomized Adversarial Training

### v2

@@ 5 Experiments
DWF introduces additional computational overhead compared to single-network baselines due to the joint training and inference of multiple experts. On our hardware, full training requires several days (approximately five days of wall-clock time, corresponding to roughly 240 GPU-hours), whereas single baselines complete substantially faster. ADVMoE and SoE exhibit intermediate training costs between these extremes. During inference, the gating module is lightweight and the dominant cost comes from evaluating the experts, resulting in approximately linear scaling with the number of experts. Although DWF is computationally heavier than plain networks, we find the overhead acceptable in practice given the considerable robustness gains. This highlights a typical robustness-efficiency trade-off.
@@
Attacks Settings Attacker’s Knowledge Defense’s Knowledge
FGSM $\epsilon\in\{0.01,~0.02,...,0.09\}$ white-box white-box
PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i\in\{10,~20,~30,~40,~50\}$ white-box white-box
Multi-Step PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i=100$, $\text{Restart}=3$ white-box black-box
Adaptive PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i=100$, $\text{EoT Samples}=1$ white-box black-box
@@
To compare our results with plain networks we used networks pretrained on CIFAR-10 from [34, 24] and networks pretrained on SVHN from [14]. These models include VGG-19 [51], ResNet-34, ResNet-50 [19], DenseNet-121, DenseNet-169 [21], GoogLeNet, Inception-V3 [53], Xception [12], and MobileNet-V2 [48].
@@
To evaluate our model’s resilience against white box attacks, we employ FGSM with the perturbation sizes $\epsilon\in\{0.01,~0.02,~\dots,~0.09\}$, and PGD with the step size $\alpha=\frac{2}{255}$, perturbation bound $\epsilon=\frac{8}{255}$, and iterations numbers $i\in\{10,~20,~30,~40,~50\}$. These setting were used during both training and test phases, highlighting that both threat model and defense system have white-box knowledge about each other. Moreover, to further evaluate our proposed defense’s performance against unseen attacks (black-box defense knowledge), we used multi-step PGD-100 with 3 restarts and adaptive PGD-100 with expectation over transformation (EoT) of one. Table 4 shows an overview of the attacks settings.
@@ Results
Table 3 illustrates the accuracy and robustness of our proposed model in comparison with other models. Figures 6 and 7 present a visual summary of the results reported in Table 3 and provide clearer comparison between our proposed model and other competitors in terms of robust accuracy under different settings for FGSM and PGD attacks. The attained results by our model are averaged over five runs, ensuring a fair comparison.
@@
6 Conclusion and Future Works
@@
While this work focuses primarily on improving robustness, training efficiency is also an important practical consideration. Fast adversarial training variants (e.g., FFAT or single-step approximations) may reduce computational cost at the expense of some robustness. A systematic study quantifying the trade-off between training time and robust accuracy for DWF under different efficient training strategies would require multiple full training runs and extensive hyperparameter tuning. We leave this investigation for future work. Moreover, We trained and tested our model against two popular attacks, FGSM and PGD. To further expand our method one can train and evaluate the model on a wider range of adversarial attacks.

### v3

@@ 5 Experiments
Baselines Trained Under Our Pipeline
ResNet-18 37.72 37.28 37.10 37.02 36.95 37.07 37
ResNet-18 (FGSM-AT) 58.65 58.21 58.17 58.14 58.13 58 58.11
ResNet-18 (PGD-AT) 58.74 58.35 58.31 58.26 58.26 58.05 58.26
@@
DWF introduces additional computational overhead compared to single-network baselines due to the joint training and inference of multiple experts. On our hardware, full training requires several days (approximately five days of wall-clock time, corresponding to roughly 240 GPU-hours), whereas single baselines complete substantially faster. ADVMoE and SoE exhibit intermediate training costs between these extremes. During inference, the gating module is lightweight and the dominant cost comes from evaluating the experts, resulting in approximately linear scaling with the number of experts.
Fractional activation evaluates all experts in parallel using continuous routing weights. As a result, inference cost scales approximately linearly with the number of experts relative to a single backbone of comparable size. The architecture does not introduce sequential routing decisions or iterative refinement steps; however, evaluating multiple experts increases FLOPs and memory proportionally to expert count. DWF is primarily a robustness-oriented design rather than a deployment-optimized solution.
@@
Attacks Settings Attacker’s Knowledge
FGSM $\epsilon\in\{0.01,~0.02,...,0.09\}$ white-box
PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i\in\{10,~20,~30,~40,~50\}$ white-box
Multi-Step PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i=100$, $\text{Restart}=3$ white-box
Adaptive PGD $\epsilon=\frac{8}{255}$, $\alpha=\frac{2}{255}$, $i=100$, $\text{EoT Samples}=1$ white-box
@@
To compare our results with plain networks we used networks pretrained on CIFAR-10 from [16] and networks pretrained on SVHN from [11]. These models include VGG-19 [36], ResNet-34, ResNet-50 [14], DenseNet-121, DenseNet-169 [15], GoogLeNet, Inception-V3 [38], Xception [9], and MobileNet-V2 [34].
@@
To evaluate our model’s resilience against white box attacks, we employ FGSM with the perturbation sizes $\epsilon\in\{0.01,~0.02,~\dots,~0.09\}$, and PGD with the step size $\alpha=\frac{2}{255}$, perturbation bound $\epsilon=\frac{8}{255}$, and iterations numbers $i\in\{10,~20,~30,~40,~50\}$. These setting were used during both training and test phases. Moreover, to further evaluate our proposed defense’s performance against unseen attacks, we used multi-step PGD-100 with 3 restarts and adaptive PGD-100 with expectation over transformation (EoT) of one. Table 4 shows an overview of the attacks settings.
@@
For each dataset, we apply a fixed per-channel normalization to all splits (train and test) using the same preprocessing pipeline. Importantly, the same normalization parameters are used consistently across all defended models, including DWF, ADVMoE, SoE, and adversarially trained ResNet-18 models, ensuring fair and controlled comparisons. The CIFAR-10 and SVHN preprocessing pipelines are dataset-specific and were not interchanged.
@@ Results
Table 3 illustrates the accuracy and robustness of our proposed model in comparison with other models. Figures 6 and 7 present a visual summary of the results reported in Table 3 and provide clearer comparison between our proposed model and other competitors in terms of robust accuracy under different settings for FGSM and PGD attacks. The attained results by our model are averaged over five runs, ensuring a fair comparison. For the defended baseline methods, we follow their standard evaluation protocols and report the corresponding single-run results.
To improve transparency and fairness, we distinguish between publicly available pretrained baselines and models trained under our experimental pipeline. Public checkpoints are evaluated using our standardized preprocessing and attack settings but retain their original training recipes. In contrast, adversarially trained ResNet-18 baselines and MoE defenses (SoE, ADVMoE, and DWF) are trained under the same optimization protocol, preprocessing pipeline, and threat model. This separation clarifies that performance differences between MoE-based methods reflect architectural and training design choices rather than inconsistencies in evaluation settings.
@@
The proposed design choices (soft routing, fractional activation, staged pretraining, and joint optimization) are not independent modifications but structurally coupled components of a unified framework. Fractional activation inherently relies on continuous routing weights, and replacing it with sparse or hard gating alters gradient flow and stability properties. Likewise, the pretrain+joint schedule is intended to promote initial expert differentiation prior to collaborative adaptation, which differs qualitatively from joint-only training. For this reason, we evaluate the framework in its complete form, as these elements collectively define the optimization behavior of the model.
6 Limitations and Future Works
@@
While this work focuses primarily on improving robustness, training efficiency is also an important practical consideration. Fast adversarial training variants (e.g., FFAT or single-step approximations) may reduce computational cost at the expense of some robustness. A systematic study quantifying the trade-off between training time and robust accuracy for DWF under different efficient training strategies would require multiple full training runs and extensive hyperparameter tuning. We leave this investigation for future work.
@@
In summary, the main practical constraints of DWF stem from computational cost, scaling behavior, and evaluation scope. Training multiple experts jointly increases training time and memory usage compared to single-network defenses, and inference cost scales approximately linearly with the number of experts, which may limit deployment in resource-constrained environments. While our evaluation focuses on strong white-box evasion attacks, robustness under broader threat models (such as transfer-based, query-limited black-box, poisoning, or physical attacks) remains to be systematically assessed.
7 Conclusion
In this paper, we introduced DWF, a soft-gated fractional mixture-of-experts defense that combines joint expert-router optimization with randomized adversarial training to improve robustness under a white-box $\ell_{\infty}$ threat model. Extensive experiments on CIFAR-10 and SVHN demonstrate that DWF consistently achieves higher robust accuracy than prior MoE-based defenses (SoE and ADVMoE) and adversarially trained single-network baselines across FGSM, multi-step PGD, and adaptive PGD attacks. Notably, these robustness gains are obtained while preserving high clean accuracy (above 90%), indicating that the proposed mixture-of-experts design mitigates the traditional robustness-accuracy trade-off.
The results show that soft gating with fractional expert activation, combined with a pretrain-then-joint training pipeline and randomized adversarial training, enables cooperative expert behavior and more stable predictions under adversarial perturbations. By isolating routing, activation, and training design within a unified MoE framework, DWF demonstrates that improving expert utilization and training dynamics can yield substantial robustness gains without requiring major architectural changes. These findings highlight the practical potential of soft-gated MoE defenses for strengthening adversarial resilience in image classification systems.